OpenStudy (anonymous):

An isotope has a half-life of 5,000 years, How long will it take to decay to 15% of its original quantity?

OpenStudy (rogue):

\[f(t) = Ae ^{kt}\]\[k = \frac {-ln 2}{t _{\frac{1}{2}}} = \frac {-ln 2}{5000} = 0.000138629436 \]We'll let A = 100, so its easy to see the percent change.\[f(t) = 100e^{0.000138629436t} = 15\]\[.15 = e^{-0.000138629436t}\]\[t = \frac {\ln .15}{-0.000138629436} \approx 13685 years\]

OpenStudy (rogue):

The constant is negative in the exponential decay equation, I made a mistake when I first wrote the equation. \[f(t) = Ae^{-kt}\]

OpenStudy (anonymous):

okay thank you. may i ask though how you found k?

OpenStudy (rogue):

Exponential decay is first order, so there is a formula that relates the half life and the constant.\[t _{\frac{1}{2}}^{} = \frac {-\ln 2}{k}\]

OpenStudy (anonymous):

okay, thank you very much!

OpenStudy (rogue):

Alright, no problem, good luck with your studies :)